OpenStudy (anonymous):

sin(theta)=3/5 solve for theta

sam (.sam.):

\[\sin(\theta)=3/5\] \[\theta=\sin^{-1}(3/5)\] \[\theta=36.87,143.13\]

OpenStudy (anonymous):

so I just plug in the inverse sin into my calculator? thank you

sam (.sam.):

yes